### v2

# Strong-field Driven Sub-cycle Band Structure Modulation and Dephasing Control

Abstract
Over the past decade, ultrafast electron dynamics in the solid state have been extensively studied using various strong light-matter interaction techniques, such as high-harmonic generation. These studies lead to multiple interpretations of light-matter interaction in the strong-field regime, with exact mechanisms not yet fully understood. It is well known that strong-field interaction with a crystalline solid leads to significant modification of its band structure and, hence, its optical properties on ultrafast timescales. In this work, we present measurements of ultrafast electric-field observables in magnesium oxide using a non-resonant nonlinear optical interaction. Using field observables, we show that strong laser fields modulate the band structure on sub-cycle timescales, thereby altering the material’s nonlinear optical response. We perform time-dependent perturbation theory calculations using a field-dependent dispersion relation and non-perturbative semiconductor Bloch equation calculations, both of which agree with experimental observations. Furthermore, we directly extract dephasing times from the real-time signal electric field envelope and show sub-cycle control of dephasing times. Our work offers a new perspective on strong-field-driven electron dynamics in solids through electric-field observables. The demonstrated attosecond modulation of the nonlinear response could have important implications for quantum light generation and quantum spectroscopy using nonlinear optical processes.
Strong-field, Ultrafast Band Modulation, Electric Field Observables, Nonlinear optics, Sub-cycle
These authors contributed equally to this work.
I Introduction
The interaction of strong, ultrafast laser pulses with solids can give rise to many intriguing phenomena that have been extensively studied using multiple ultrafast techniques. Examples include high-harmonic generation (HHG) [16, 17], the realization of the Haldane model in a laser-dressed crystal [30], and transient optical response modulation, also known as the dynamical Franz-Keldysh effect [22, 34, 28, 39]. These effects are fundamental to understanding electron dynamics in the strong field regime. Several theoretical models have been developed to study these phenomena, such as the Houston basis formalism [55, 59] and Wannier functions [36, 48], and the Semiconductor Bloch equation [43, 20, 21] to model HHG in solids. To understand the dynamical Franz-Keldysh effect, more sophisticated time-dependent density functional theory (TDDFT) techniques have been used [28, 37]. Despite differences in approach, these models inherently incorporate non-perturbative strong-field effects in the material to gain insight into ultrafast electron dynamics, as measured by various experimental observables.
In the study presented here, we investigate the effects of electron interactions with strong fields in solids using ultrafast nonlinear optical spectroscopy, specifically electric-field observables in four-wave mixing (FWM). This nonlinear spectroscopy technique involves three coherent femtosecond pulses interacting with a crystalline solid to produce a fourth pulse that is proportional to the material’s third-order susceptibility ($\chi^{(3)}$) [4]. Previously, the FWM process has been studied in quantum well and quantum dot systems experimentally [29, 33] and theoretically [11, 14]. However, in an optically thick semiconductor, the propagation effect of the FWM output pulse becomes complex [61, 44, 41]. Recently, with advances in Free-Electron Lasers (FELs), the FWM process using purely extreme ultraviolet (XUV) pulses has also been demonstrated [2, 15]. On the other hand, nonlinear response functions have also been measured using HHG in solids [19].
In this article, we report an experiment in which we measure electric-field observables for degenerate four-wave mixing (DFWM) in magnesium oxide (MgO) to study the strong-field modification of the band structure and, hence, the nonlinear response function in the material. Using the TADPOLE technique [13], the amplitude and phase of the DFWM output field are measured, providing complete information about ultrafast nonlinear light-matter interactions. We also develop a theoretical model based on time-dependent perturbation theory, combined with a field-dependent dispersion relation [18], to simulate the DFWM process in a large-band-gap crystal (such as MgO). We study the modulation of the amplitude and the temporal chirp of the output signal as we vary the time delay between incident pulses with attosecond resolution. It was found that both the amplitude and the chirp oscillate with a time delay, and our model, with good agreement, predicts that this oscillation can occur only due to ultrafast modulation of the band structure of MgO under strong electric fields. Additionally, we directly extract dephasing times from our measured signal electric field and show sub-cycle control of dephasing time due to the strong field driven modulation of the band structure.
Figure 1: (a) Ultrafast Band modulation: The net electric field resulting from the superposition of three ultrafast femtosecond pulses interacts with MgO and significantly modifiesthe band structure of the material. This interaction modulates the band structure (represented by opaque curves) and shifts it (transparent curves) from its original position on an ultrafast timescale. The measured output electric field from the nonlinear process can capture this modulation, providing insights into the dynamics of strong-field interactions in a solid. (b) Schematic diagram of the spectral interferometry measurement for degenerate four-wave mixing (DFWM): Experimental setup illustrates the interaction of three ultrafast pulses (Gate 1, Gate 2, and Probe) with the magnesium oxide crystal and the generation of a DFWM signal pulse. This signal interferes with a reference pulse at a spectrometer, resulting in a spectral interference measurement. R - Reference pulse, S - Signal pulse, and BS - beam splitter.
Figure 2: Experimentally Measured Amplitude and Phase: (a) Amplitude $|E_{0}(t,\tau)|$ and (b) Phase $\varphi(t,\tau)$ of the Degenerate Four-Wave Mixing (DFWM) electric field signal in Magnesium Oxide (MgO). (c) Time integrated temporal amplitude and (d) Temporal Chirp as a function of Time Delay $\tau$ with a $\tau$ delay step of 200 as extracted from (a) and (b), respectively. The shaded region represents the statistical error in the measurement.
II Results and Discussion
II.1 Amplitude and phase of the measured DFWM electric field
Our experiment employs the spectral interferometry-based technique of TADPOLE [13] to completely characterize the degenerate Four-Wave Mixing (DFWM) signal generated using three 60 fs near-infrared (NIR) pulses centered around 800 nm ($\hbar\omega_{0}\approx 1.56eV$). These three linearly polarized pulses interact with the Magnesium Oxide (MgO) crystal with an intensity of $\sim 0.1$ TW/cm2 in a BOXCAR configuration [47], producing a fourth beam in a new direction that satisfies the phase matching condition $\bm{k_{s}}=\bm{k}_{1}-\bm{k}_{2}+\bm{k}_{3}$. Figure 1 (a) illustrates the schematic representation of the highest valence band and the lowest conduction band of MgO, along with the Four-Wave Mixing (FWM) process occurring within the crystal. This FWM process occurs with a simultaneous modulation of the band by the net electric field present during the interaction, as discussed later. Since the MgO bandgap is about 7.8 eV and the photon energy of each incoming pulse is 1.56 eV, the interaction is highly off-resonant. A reference beam then interferes with the output signal, and spectral interference is measured using a spectrometer. Representative interference fringes from the measurement along with the experimental setup are shown in Figure 1 (b). From this interference, we can extract the temporal amplitude and phase of the electric field of the DFWM signal (more details in Section IV.1). The electric field is then measured for each time delay ($\tau$) between two phase-locked Gate pulses and the Probe pulse, with a time delay resolution of 200 as. Therefore, the complete electric field of the DFWM signal can be expressed as $E(t,\tau)=|E_{0}(t,\tau)|\exp{\left(-i\varphi(t,\tau)\right)}$, where the temporal amplitude $|E_{0}(t,\tau)|$ and phase $\varphi(t,\tau)$ depend on real time $t$ and the time delay between the Gate and Probe pulses $\tau$. Both the amplitude and the phase are plotted in Figures 2 (a) and (b), respectively. For each time delay $\tau$, the phase is fit to a fifth-order polynomial,
$\varphi(t,\tau)=\varphi_{0}+a(\tau)t+b(\tau)t^{2}+c(\tau)t^{3}+d(\tau)t^{4}+e(\tau)t^{5}$ (1)
The second-order coefficient $b(\tau)$, which is also called the temporal chirp, is extracted as a function of the time delay $\tau$. We integrate the extracted amplitude over the total pulse duration for each time delay; this is called the time-integrated amplitude and is defined as $\int|E_{0}(t,\tau)|\ dt$. The time-integrated amplitude as a function of time delay is shown in Figure 2 (c). Figure 2 (d) shows the chirp parameter $b(\tau)$ as a function of the time delay with attosecond resolution. Modulation of nonlinear signals has previously been observed in semiconductors, where a strong near-infrared (NIR) pump pulse generates mid-infrared (MIR) sidebands by mixing with a strong THz pulse [9]. The time delay between the two mixing pulses modulates the spectral intensity of the MIR sideband. However, sub-cycle modulation in electric field observables, as shown here, has not been observed previously.
II.2 Strong-field Driven Nonlinear Response Modulation
The output DFWM signal depends on the third-order polarization of the crystal. The spectrum of third-order polarization can be written in terms of the nonlinear response function, i.e., third-order susceptibility $\chi^{(3)}(\omega;\omega,-\omega,\omega)$, which is a material property. However, in strong fields, higher-order contributions can arise. Most generally, the polarization is [4],
$P_{i}(\omega)=\chi^{(3)}_{ijkl}\mathcal{E}_{j}(\omega)\mathcal{E}_{k}(\omega)\mathcal{E}_{l}(\omega)+\chi^{(5)}_{ijklmn}\mathcal{E}_{j}(\omega)\mathcal{E}_{k}(\omega)\mathcal{E}_{l}(\omega)\mathcal{E}_{m}(\omega)\mathcal{E}_{n}(\omega)+\cdots$ (2)
If the polarization vectors of the input and output electric fields are in the same direction, as in our experiment, the indices $i,j,k,\cdots$ are the same. For the remainder of this discussion, we omit the polarization subscripts for brevity. Here, $\mathcal{E}(\omega)=(E_{1}(\omega)e^{i\bm{k_{1}}\cdot r}+E_{2}(\omega)e^{i\bm{k_{2}}\cdot r}+E_{3}(\omega)e^{i\bm{k_{3}}\cdot r}+\mathrm{c.c.})$. Based on the phase-matching condition mentioned above, the only terms that contribute to the measured signal are
$P_{DFWM}=$ $\chi^{(3)}_{0}E_{1}(\omega)E_{2}^{*}(\omega)E_{3}(\omega)e^{i(\bm{k}_{1}-\bm{k}_{2}+\bm{k}_{3})\cdot\bm{r}}$
$+\chi^{(5)}_{0}E_{1}(\omega)E_{2}^{*}(\omega)E_{3}(\omega)(|E_{1}|^{2}+|E_{2}|^{2}+|E_{3}|^{2})e^{i(\bm{k}_{1}-\bm{k}_{2}+\bm{k}_{3})\cdot\bm{r}}+\cdots$ (3)
Here, $E_{1},E_{2}$ and $E_{3}$ are the two Gate and Probe pulse electric fields in the frequency domain. The introduction of a time delay $\tau$ between Gate and Probe pulses will result in a phase term of $\exp{(-i\omega\tau)}$ in the electric field of the Gate pulses. However, this phase term does not result in a $\tau$-dependent modulation of the amplitude or phase of the polarization corresponding to the DFWM process when $\chi_{0}^{(3)}$ is independent of the input fields (for a detailed calculation, see the supplementary material). In the presence of moderately strong fields, nonlinear response functions are expected to be significantly modified as observed, for example, in studies of the dynamical Franz-Keldysh effect using attosecond transient absorption spectroscopy [28, 9]. Thus, we model the attosecond modulation of the nonlinear response function due to the strong field interaction of electrons in the crystal such that $\chi^{(3)}\rightarrow\chi^{(3)}(\mathcal{E})$ becomes an instantaneous function of the incoming electric fields.
In high-intensity laser fields, excited electron-hole pairs can extend beyond the edge of the Brillouin zone, leading to Landau–Dykhne–type transitions. Such effects have been observed in high-harmonic spectra [52, 53]. These transitions provide a powerful means to probe modifications of the band structure induced by intense laser pulses. To directly probe band-structure modifications, time- and angle-resolved photoelectron spectroscopy (Tr-ARPES) can be employed, as it images electron dynamics in materials with simultaneous momentum and energy resolution. Ab initio simulations have been performed to study Tr-ARPES in periodic systems under strong fields [32]. Most recently, a Tr-ARPES experiment on bismuth telluride observed field-dressed bands in the presence of an intense laser field [57]. This suggests that the ultrafast modification of the band structure (or adiabatic field-dressed bands) offers a useful framework for understanding off-resonant electron interactions in strong-field regimes. Alternative perspectives for interpreting these field-dressed bands have also been proposed, most notably within the framework of Floquet theory [50].
For a rigorous quantitative analysis, we start with time-dependent perturbation theory to simulate the DFWM experiment. However, the time-integrated amplitude and the chirp of the calculated signal electric field do not exhibit any of the oscillatory features observed in the experiment (see supplementary material figure 8 red vs blue curves). This aligns with the expectation that the field-free response function accounts only for perturbative wave-mixing processes [4, 8] and cannot include any field-driven modification of the response function. We thus introduce an ultrafast band-modulation model [18] to implicitly account for the strong-field effects on the nonlinear response function. In the presence of an external field, the crystal momentum of the electrons can be written as $\mathbf{p}(t)=\mathbf{p}_{0}+e\mathbf{A}_{ext}(t)$ [55, 24, 23, 17, 18, 59], where $\bm{p}_{0}$ is the initial momentum of the electron and $\bm{A}_{ext}(t)$ is the vector potential of the external field. This results in the time-dependent modification of the electronic dispersion relation, i.e. $\epsilon(\mathbf{p})\rightarrow\epsilon(\mathbf{p}_{0}+e\mathbf{A}_{ext}(t))$ where $\epsilon$ is the energy of the electron. The details of the calculations are listed in the supplementary material. A theoretical investigation of the nonlinear optical response under strong light-matter interaction in MgO has previously been carried out using the Floquet–Bloch formalism [38]. Here, we demonstrate that incorporating non-perturbative effects (specifically, band-structure modulation) within a perturbative framework yields excellent agreement with the experiment.
Figure 3: Fourier Transformation in Time Delay: The Fourier transforms ($\mathcal{FT}$) with respect to the time delay of the experimentally measured (a) time-integrated amplitude and (b) temporal chirp parameter are shown. These results are compared with the corresponding (c) time-integrated amplitude and (d) chirp parameter obtained from the electric field calculated using time-dependent (TD) perturbation theory, which includes the band-structure modulation induced by the strong field (see the main text for details). The calculation parameters are: IR pulse intensity $I_{0}=0.2\ \mathrm{TW/cm^{2}}$, photon energy $\hbar\omega_{0}=1.56\ \mathrm{eV}$, and MgO bandgap $\Delta=7.78\ \mathrm{eV}$. The (e) time-integrated amplitude and (f) temporal chirp are also extracted from the calculated output electric field using the Semiconductor Bloch Equations (SBE), where the contributions from intraband and interband currents are shown separately.
We perform a Fourier transformation ($\mathcal{FT}$) of our field observables in the time-delay domain. Figure 3 (a) and (b) show the $\mathcal{FT}$ of the time-integrated amplitude and the chirp of the measured DFWM electric field shown in Figure 2, respectively. We emphasize that the Fourier transformation is in the time-delay $\tau$ domain and not in real time $t$, and the observed frequency components reflect delay-dependent modulations. We compare these with the theoretically calculated field observables after the inclusion of ultrafast modulation of the band structure within time-dependent perturbation theory in Fig. 3 (c) and (d). The calculation shows good agreement with the experimentally measured field observables for an input intensity of 0.2 TW/cm2, which is close to the experimentally measured intensity of $\sim$ 0.1 TW/cm2. The uncertainty in experimentally measured intensity is typically large for femtosecond pulses due to a complex beam profile at the focal spot. The time-integrated amplitude exhibits a strong $1\omega_{0}$ component of the delay frequency ($\omega_{0}$ is the central frequency of the NIR pulse). This is because the amplitude of the net incoming electric field modulates with the time delay between the Gate and the Probe pulses. As a result, the nonlinear response function is modulated via adiabatic field-dressed bands (see supplementary material section A for details). The chirp, as shown in Figure 3 (b) and (d), is more sensitive to such modulation of the nonlinear response function. Unlike the time-integrated amplitude, the $\mathcal{FT}$ of the chirp shows a significant contribution from higher frequency components ($1\omega_{0}$).
To compare our results to a non-perturbative simulation of our experiment, we solve the semiconductor Bloch equations (SBE) for MgO with our experimental parameters as inputs [43, 20, 21]. Unlike perturbative models, the non-perturbative SBE framework can separate contributions from intra-band and inter-band currents. The time-delay Fourier transforms of the time-integrated amplitude and the temporal chirp, corresponding to the electric fields generated by intra- and inter-band currents, are shown in Fig. 3 (e) and (f). These plots show that the interband and intraband contributions to the DFWM signal-field observables are comparable under our experimental conditions. Further details of the calculation are provided in the supplementary material.
[…]
The simple model of nonlinear susceptibility modulation discussed earlier provides physical insight into the emergence of peaks at integer multiples of $\omega_{0}$ observed in both the experimentally measured and theoretically calculated field observables in Fig. 3. We start from Eq. 2, where the third-order susceptibility ($\chi^{(3)}(\mathcal{E})$) is written as an instantaneous function of the total interacting electric field $\mathcal{E}$. The total electric field $\mathcal{E}$ is the sum of the two gate pulses ($E_{1}$, $E_{2}$) and the probe pulse ($E_{3}$) electric field. With an introduction of time delay ($\tau$) between $E_{1}$ (phase-locked with $E_{2}$) and $E_{3}$, the three pulses $E_{1},E_{2}$ and $E_{3}$ in the frequency domain can be written as,
$E_{1}(t-\tau)$ $\leftrightarrow E_{1}(\omega)e^{i(\bm{k}_{1}\cdot\bm{r}-\omega\tau)}+c.c$ (4a)
$E_{2}(t-\tau)$ $\leftrightarrow E_{2}(\omega)e^{i(\bm{k}_{2}\cdot\bm{r}-\omega\tau)}+c.c$ (4b)
$E_{3}(t)$ $\leftrightarrow E_{3}(\omega)e^{i\bm{k}_{3}\cdot\bm{r}}+c.c.$ (4c)
where $\bm{k}_{i}$ is the wave vector of the $i^{th}$ field. We also expand the modified nonlinear susceptibility in the power series of $\mathcal{E}$ (net electric field) as $\chi^{(3)}(\mathcal{E})=\chi^{(3)}_{0}+a\mathcal{E}+b\mathcal{E}^{2}+\mathcal{O}(\mathcal{E}^{3})$, where $\chi^{(3)}_{0}$ is the field-free third order susceptibility. The parameters $a$ and $b$ represent the coefficients governing the first- and second-order modulations of the nonlinear susceptibility. Plugging in all the relations, we get the expression of an effective DFWM polarization as,
$P^{(3)}_{eff}$= $\left[\chi^{(3)}_{0}+a\left(E_{1}(\omega)e^{-i\omega\tau}+E_{2}(\omega)e^{-i\omega\tau}+E_{3}(\omega)+c.c.\right)\right$.
$+\left.b\left(E_{1}(\omega)e^{-i\omega\tau}+E_{2}(\omega)e^{-i\omega\tau}+E_{3}(\omega)+c.c.\right)^{2}+\cdots\right]E_{1}(\omega)E_{2}^{*}(\omega)E_{3}(\omega)e^{i(\bm{k}_{1}-\bm{k}_{2}+\bm{k}_{3})\cdot\bm{r}}.$ (5)
Here, it is important to note that the phase-matching conditions arising from terms that include the wave vectors $\bm{k_{i}}$ corresponding to the input fields only apply to the wave-mixing terms outside the brackets. We do not include wave-vector terms in the field-dependent expansion of $\chi^{(3)}$ because the field-driven modification of $\chi^{(3)}$ is an instantaneous field effect independent of the direction of propagation of the field(s). This allows us to model non-perturbative field modification of the response function within a third-order perturbative model and is distinct from a higher-order perturbation expansion shown in Eq. II.2. Using this expression, we can also derive an analytical expression for the temporal chirp (see the supplementary material for a detailed derivation).
Figure 5: Extraction of dephasing times: (a) Shows a schematic of DFWM in MgO, where the asymmetry in the output pulse contains the information of third-order coherence dephasing. The asymmetry shown in the output pulse is exaggerated for illustrative purposes. (b) A representative plot of the real-time amplitude of the measured output electric field at $\tau=26$ fs fitted with an exponentially modified Gaussian (EMG) function (black, solid curve) is shown. See supplementary material for details. (c) Dephasing times as a function of time delay are extracted, showing that strong light-matter interaction leads to attosecond modulation and control of dephasing. The shaded region represents the fitting error.
It can be readily seen from Eq. II.2 that the $1\omega_{0}$ peak in Fig. 3 primarily originates from the linear expansion of the net field-dependent $\chi^{(3)}$. The second-order expansion contributes not only to the $1\omega_{0}$ peak but also generates a component at $2\omega_{0}$. A similar reasoning extends naturally to the higher-order expansions of $\chi^{(3)}$. Within this framework, the difference in the relative peak strengths at $1\omega_{0}$ and $2\omega_{0}$ can be understood by noting that the third-order susceptibility is predominantly governed by the linear term, i.e., $\chi^{(3)}\approx\chi_{0}^{(3)}+a\mathcal{E}$. Analogous phenomena have previously been observed in which an external field modifies both the linear and nonlinear response functions of a material. A well-known example is the optical Kerr effect, in which the change in the linear refractive index of a medium is proportional to the intensity of the applied field [49, 42, 5]. Similar effects have also been utilized in electric-field-induced second-harmonic generation, where the external field breaks the inversion symmetry of otherwise centrosymmetric systems, making the second-order response function nonzero[58, 60]. However, in this experiment, the observed oscillation cannot be explained by including a fifth-order term under the phase-matching condition, as shown previously using Eq. II.2. To further support this interpretation, we isolate the $1\omega_{0}$ peak of the measured chirp in Fig. 3 (b) and inverse Fourier transform to obtain the single frequency filtered oscillation of the temporal chirp, shown in Fig. 4 (a). This is then compared with the $1\omega_{0}$ component of the chirp calculated using Eq. II.2. Fig. 4 (b) shows the time delay-dependent phase of the oscillations in Fig. 4 (a). The model accurately reproduces the $\pi$-phase jump in the chirp oscillation near $\tau=20~\mathrm{fs}$, in excellent agreement with the experimental observation, as shown in Fig. 4 (b). The offset of this phase jump from zero time delay is due to the chirp present in the input fields.
II.3 Sub-cycle control of dephasing
From perturbation theory, the third-order polarization can be written as[4, 31],
$P^{(3)}(t)$ $=\operatorname{Tr}\left[\hat{\mu}\hat{\rho}^{(3)}(t)e^{-t/\tau_{\text{dephase}}}\right]$ (6)
In the density matrix formalism, the diagonal terms correspond to the population of electronic states, and the off-diagonal elements represent coherence between states. As coherences evolve in time phase fluctuations attenuate their temporal evolution, a process called dephasing [31]. To model dephasing of coherences, phenomenological dephasing terms $\gamma_{\text{dephase}}=1/\tau_{\text{dephase}}$ are added. This means that the output-field pulse width depends on the dephasing of the coherence created by the probe pulse in the final step of the nonlinear interaction. To quantify third-order coherence dephasing, we fit the measured output temporal field with an Exponentially Modified Gaussian (EMG) function. We fix the Gaussian width ($\sigma$) to 21.2 fs (corresponding to the input probe pulse FWHM of 50 fs) and extract the dephasing time $\tau$ of the DFWM signal field as shown in Fig. 5 (b) (see supplementary materials for details on fitting). The extracted dephasing times $\tau_{\text{dephase}}$ from the EMG fit are plotted as a function of time delay in Fig. 5 (c). The data indicate that band-structure modification resulting from strong light-matter interaction modulates the dephasing times of nonlinear interactions on sub-cycle time scales. Multiple factors could influence dephasing times in solids, including doping, electron-electron interactions, and electron-phonon interactions. In this study, the measured dephasing time likely reflects electron-electron interactions that affect real-space coherence of electron-hole pairs across lattice sites, as in the case of HHG [7]. Furthermore, band-structure modulation induced by strong light-matter interaction also alters electron-electron interactions, thereby modulating dephasing times.
III Conclusion
In summary, we have measured the complete electric field of the degenerate four-wave mixing (DFWM) signal in magnesium oxide using spectral interferometry. By analyzing field observables, such as the time-integrated amplitude and the signal field chirp, as functions of the sub-cycle time delay between incoming pulses, we identified unique oscillatory features. We attribute these features to the ultrafast modulation of the nonlinear response function of the crystal due to modulation of the band structure. To support this interpretation, we simulated the DFWM experiment using time-dependent perturbation theory and found that oscillatory behavior could be reproduced only when instantaneous ultrafast band-structure modulation driven by the strong field is included. Such ultrafast band modulation effects are known to leave signatures in other optical observables, such as high-harmonic spectra [52, 53, 25]. Here, we demonstrate that the influence of field-dressed bands is clearly manifested in the field observables of nonlinear optical signals. Our results indicate that field observables can serve as direct probes of strong-field-driven attosecond electron dynamics in solid-state systems. We have demonstrated that modeling of our experimental observations requires incorporating strong light–matter interactions into the non-perturbative modification of the third-order nonlinear susceptibility. Furthermore, by fitting the real-time electric field envelope of the signal pulses, we have extracted dephasing times corresponding to this non-resonant nonlinear interaction. This represents a direct measurement of dephasing in a nonlinear process via electric field measurement opening doors to enhacing our understanding of dephasing mechanisms. The modulation of the band structure is seen to modulate the dephasing time, resulting in sub-cycle control of dephasing.
Our findings have significant implications for the manipulation and control of observables governed by the nonlinear response function. For instance, it is well established that signals generated through nonlinear optical processes can serve as sources of non-classical light. Processes such as parametric down-conversion (PDC) and degenerate four-wave mixing (DFWM) are known to produce squeezed light [3, 6, 1, 35, 12]. The degree of squeezing in these processes is directly proportional to the nonlinear coupling coefficients of the respective systems, i.e., $\propto\chi^{(2)}$ for PDC and $\propto\chi^{(3)}$ for DFWM [12]. It has been demonstrated that squeezing in DFWM signals can also be observed from ultrafast light–matter interactions [40, 45, 62]. With the strong sub-cycle modulation of $\chi^{(3)}$ demonstrated in this work, it is now possible to control the degree of squeezing on attosecond time scales [62]. This could pave the path to achieve extreme levels of squeezing based on field induced nonlinearities that exceed current squeezing levels achieved using periodically-poled nonlinear materials [54] with numerous applications in quantum metrology, quantum spectroscopy, and quantum information science.
IV Methods
IV.1 Experimental Setup
In our setup of the DFWM experiment, as seen in Fig. 1, three pulses with different propagation vectors interact with the medium in a non-collinear geometry and generate a fourth pulse due to a third-order nonlinear interaction in a BOXCAR configuration. Due to the non-collinear geometry, the signal pulse is spatially separated from the other three beams. The output signal has a momentum of $\bm{k_{s}}=\bm{k}_{1}-\bm{k}_{2}+\bm{k}_{3}$. The Probe and Gate beams are centered at 800 nm, with a pulse width of 50 fs and a repetition rate of 1 kHz. The Gate pulse is divided into two phase-locked Gate pulses (Gate 1 and Gate 2) using a mask. The Probe pulse is delayed with respect to the Gate pulses using an optical delay stage, and the Gate polarization can be rotated relative to the Probe polarization. However, in this experiment, the polarizations of all four pulses are fixed along the same direction. The MgO crystal is mounted on a controllable rotation mount to allow precise control of the crystal angle relative to the probe polarization. The Probe and the two Gate pulses each have a pulse energy of $\sim$ 500 nJ and form the input pulses in our DFWM measurement. Before the interaction, a Reference pulse is derived from the Probe pulse. After the interaction, the Signal pulse is isolated by an iris, so the residual Probe and Gate beams are blocked. The Signal pulse passes through a high-contrast polarizer, which extracts the polarization component of the signal along the input polarization direction. The output signal pulse was completely measured using TADPOLE [13], where the Reference pulse is combined with the Signal pulse in a spectrometer to obtain the measurement of spectral interference. Frequency-Resolved Optical Gating (FROG) [51] is used to completely characterize the Reference pulse. The measured spectral interference of the DFWM signal and the Reference can be written as:
$\begin{split}S(\omega)&=S_{R}(\omega)+S_{S}(\omega)+\sqrt{S_{R}(\omega)}\sqrt{S_{S}(\omega)}\cos(\varphi_{S}(\omega)-\varphi_{R}(\omega)+\omega\,\tau_{R})\end{split}$ (7)
$S(\omega)$ is the spectral intensity, $\omega$ is the angular frequency, and $\varphi(\omega)$ is the spectral phase. The subscripts $S$ and $R$ stand for signal and reference, respectively. $\tau_{R}$ is the time delay between the Reference and the Signal and the Reference and the background.
IV.2 Simulation
To simulate the complete electric field of the DFWM signal, we employ time-dependent perturbation theory to evaluate the third-order expectation value of the microscopic current operator. The streaking of the electron momentum, $\bm{p}(t)=\bm{p}0+e\bm{A}_{ext}(t)$, is explicitly incorporated. The emitted DFWM signal is expressed as $\bm{E}_{\mathrm{DFWM}}(t)\propto\frac{\partial}{\partial t}\langle\bm{J}(t)\rangle^{(3)}$, where
$\langle J^{i}(t)\rangle^{(3)}=$ $\sum_{\alpha,\beta,\gamma,\bm{q}\in\mathrm{BZ}}\mathcal{N}^{i\alpha\beta\gamma}_{\mathbf{q}}\int_{t_{0}}^{t}dt_{3}\int_{t_{0}}^{t_{3}}dt_{2}\int_{t_{0}}^{t_{2}}dt_{1}\ A_{ext}^{\alpha}(t_{1})A_{ext}^{\beta}(t_{2})A_{ext}^{\gamma}(t_{3})$ (8)
$\times\exp\left(i\int_{t_{0}}^{t}\Theta_{0,1,\mathbf{q}}(t^{\prime})dt^{\prime}+i\int_{t_{0}}^{t_{1}}\Theta_{1,0,\mathbf{q}}(t^{\prime})dt^{\prime}\right.+\left.i\int_{t_{0}}^{t_{2}}\Theta_{0,1,\mathbf{q}}(t^{\prime})dt^{\prime}+i\int_{t_{0}}^{t_{3}}\Theta_{1,0,\mathbf{q}}(t^{\prime})dt^{\prime}\right).$
Here $\Theta_{1,0,\mathbf{q}}(t)$ is the time-dependent energy gap, defined as $\Theta_{1,0,\mathbf{q}}(t)=\epsilon_{\mathrm{CB}}(t)-\epsilon_{\mathrm{VB}}(t)$. Here, 0 and $1$ denote the valence band (VB) and conduction band (CB), respectively, in the two-band model, and $i,\alpha,\beta,\gamma$ index the polarization directions. Further details of the derivation are provided in the Supplementary Material.
Acknowledgments
This work was primarily supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award #\# DE-SC0024234. NS acknowledges support from the National Science Foundation under award #\# 2208061, which supported YZ. Work at the Molecular Foundry was supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility (project m4269). The authors thank Dr. Rudro Biswas for helpful discussions.